This issue should now be resolved.
We are currently looking into an issue on this server
13th February 2006, at 7:24AM
13th February 2006, at 7:29AM
We are currently looking into an issue affecting the business hosting server.
This issue has now been resolved, and the server is serving sites again.
11th February 2006, at 3:13PM
11th February 2006, at 3:27PM
This issue has now been resolved.
We are currently looking into an issue affecting the sonny.34SP.com server
5th February 2006, at 2:02PM
5th February 2006, at 2:04PM
We are presently investigating a failure of the mysql server on the server at this time.
This issue has now been resolved
1st February 2006, at 9:48PM
1st February 2006, at 9:54PM
We are currently experiencing problems with our core network. This problem is being investigated at this time and we hope to resolve the issue as soon as possible.
We are currently liasing with the technical team at our connectivity providers and are continuing to investigate the cause of this problem.
Network access has been restored - we are still looking into the causes of this with a view to preventing similar occurrences in the future.
The problem has re-occurred. At this time it looks as though we are being subject to a DoS attack and are doing all that we can to mitigate its effects.
The attack was blocked and all is now running as it should. We will be keeping a close eye on the situation.
29th January 2006, at 3:06PM
29th January 2006, at 10:10PM
We are presently upgrading core elements of the 34SP.com network. You may find your services unavailable or slowed temporarily.
The upgrade has now been completed. Services are now fully restored.
27th January 2006, at 3:28AM
27th January 2006, at 4:49AM
We are currently investigating an issue with plesk on the bishop server. This will also impact on database services.
The issue was related to an incorrect restart of the database server and has now been resolved.
24th January 2006, at 11:29PM
24th January 2006, at 11:50PM
server3.xenserve.com is presently refusing connections.
The machine has been power cycled and we are now waiting on the results from this.
The machine has now been restored.
24th January 2006, at 4:42PM
24th January 2006, at 4:48PM
An issue on the oversoul server is currently being looked into.
This matter is being looked into further after a repeat.
Steps have been taken to stop the problem on the server. We are monitoring the situation but the matter should be resolved.
20th January 2006, at 9:12AM
20th January 2006, at 10:35AM
Business1.34sp.com is now fully restored.
The business1.34sp.com server has been rebooted due to failure.
The server has failed to recover from the reboot and an engineer has been dispatched to the data center to investigate the nature of the failure.
We anticipate their arrival within 60 minutes.
18th January 2006, at 3:34PM
18th January 2006, at 5:40PM